OpenStudy (anonymous):

4 identical hollow cylindrical columns of steel support a big structure of mass 50,000 kg.the inner and outer radii of the columns are 30cm and 40 cm respectively . Assuming load distribution to be uniform,calculate the compressional strain of each column. what formula should i use for this? @IIT study group

OpenStudy (anonymous):

oh I have to revise this section! elasticity and stuff, sorry

OpenStudy (anonymous):

its okie...:( u know youngs modulus n things lyk that??

OpenStudy (aravindg):

Young’s modulus of steel, Y = 2 × 1011 Pa

OpenStudy (aravindg):

Total force exerted, F = Mg = 50000 × 9.8 N

OpenStudy (aravindg):

Stress = Force exerted on a single column =(50000*9.8)/4 =122500 N

OpenStudy (aravindg):

Y= stress/strain

OpenStudy (aravindg):

strain=FY/A

OpenStudy (aravindg):

SRRY

OpenStudy (aravindg):

strain =stress/Y

OpenStudy (aravindg):

HERE STRESS =122500

OpenStudy (aravindg):

Oh

OpenStudy (aravindg):

got it

OpenStudy (aravindg):

strain=F/AY

OpenStudy (aravindg):

STRAIN= (122500)/[3.14*((.6)^2-(.3)^2))*2*10^11]

OpenStudy (aravindg):

Hence, the compressional strain of each column is 7.22 × 10^–7

OpenStudy (aravindg):

oh hw nicely done:)

OpenStudy (anonymous):

tnank youuuu... :)

OpenStudy (aravindg):

wlcm